# Hear-Your-Click: Interactive Video-to-Audio Generation via Object-aware Contrastive Audio-Visual Fine-tuning

### v1

@@ 4.1.1. Dataset
Our models are trained and evaluated on the VGG-AnimSeg dataset, which consists of 27,200 training samples and 2,720 testing samples. When data augmentation via Random Visual Sampling (RVS) is applied, an additional 6,800 augmented training samples are incorporated, resulting in a total of 34,000 samples for model training.
@@ 4.1.2. Evaluation Metrics
For quantitative assessment, we use Frechet Distance (FD), Frechet Audio Distance (FAD), Inception Score (IS), Kullback-Leibler Divergence (KL) and Kernel Inception Distance (KID) following other audio generation studies (Iashin and Rahtu, 2021; Liu et al., 2023, 2024; Luo et al., 2024; Xing et al., 2024). FD, FAD, KL and KID are employed to quantify the similarity between the generated audio and the ground truth, while IS is utilized to assess the diversity and quality of the generated audio samples. To better evaluate the correspondence between video and audio elements, we introduce a new metric, the CAV score. The score leverages the model from C-MCR (Wang et al., 2023), which integrates CLIP (Radford et al., 2021) and CLAP (Elizalde et al., 2023) to provide audio-image contrastive representations. For each audio-video pair, we obtain per-frame image embeddings and an audio embedding, and then calculate the similarity between the average image embedding and the audio embedding. A higher CAV score indicates a better match between the objects in the generated audio and the original video.
@@
Figure 6. Visualization of the generated samples. (a) Our method successfully reproduces similar sounds with the ground truth while preserving synchronization, demonstrating great performance in capturing transient motion. (b) As the cows run away, the volume of the audio generated by our method diminishes, while preserving finer details.
@@ 4.1.3. Implementation Details
For the VOS setting, the IoU threshold and NMS threshold of SAM are set to 0.88 and 0.8. TAM is configured to use 15 voting frames. For data preprocessing, we resample 10-second video clips to 4 fps and resize each frame to 224$\times$224 pixels. For audio samples, we compute Mel spectrograms with a hop size of 250 and 128 mel bins during OCAV training, and a hop size of 256 and 128 mel bins during LDM training. Considering the limited size of our dataset, which is insufficient for training the model from scratch, we construct our framework based on Diff-Foley and fine-tune our model using its pre-trained weights. Specifically, the $f_{v}$ and $f_{m}$ in MVE utilize the SlowOnly (Feichtenhofer et al., 2019) architecture, known for its proficiency in detecting temporal action changes within videos. $f_{a}$ leverages an audio encoder provided by PANNs (Kong et al., 2020). The architecture of LDM, including the latent encoder and decoder, is inherited from Stable Diffusion-V1.4 (SD-V1.4) (Rombach et al., 2022). During the OCAV training process, we randomly sample 4-second segments from the original audio and videos, denoted as $\mathcal{A}\in\mathbb{R}^{256\times 128}$ and $\mathcal{V}\in\mathbb{R}^{16\times 224\times 224\times 3}$, yielding corresponding features $\boldsymbol{x}_{a}\in\mathbb{R}^{16\times 512}$, $\boldsymbol{x}_{v}\in\mathbb{R}^{16\times 512}$. For LDM training, we first extract visual features through MVE from the 10-second videos, resulting in $\boldsymbol{x}_{v}\in\mathbb{R}^{40\times 512}$. Following the Im2Wav (Sheffer and Adi, 2023) method, we extract per-frame CLIP embeddings, resulting in $\boldsymbol{x}^{*}_{v}\in\mathbb{R}^{40\times 512}$, and add them to $\boldsymbol{x}_{v}$ to form the condition embedding $\boldsymbol{x}_{c}$ for the LDM. At the inference stage, we set the Classifier-Free Guidance Scale to 4.5, the Classifier-Guidance Scale to 50, and use the DPM-Solver (Lu et al., 2022) sampler with 50 inference steps.
@@ 4.2.1. Qualitative Results
We conduct a comparative experiment of our method with several recently published, open-source V2A approaches, including SpecVQGAN (Iashin and Rahtu, 2021), Im2Wav (Sheffer and Adi, 2023), Seeing and Hearing (SAH) (Xing et al., 2024) and Diff-Foley (Luo et al., 2024). The qualitative evaluation focuses on three main aspects: the capacity to produce coherent audio, the relevance of the generated audio to the target objects, and the temporal alignment between audio and visual content, particularly focusing on the movements within it. As shown in Fig. 6, our method excels in generating audio most similar to the original for given video inputs, accurately capturing nuances such dog barks and cow moos, while maintaining synchronization with their movements. Other methods fall short, either producing visually accurate but poorly synchronized or low-quality audio (Im2Wav), or generating irrelevant sounds (SAH). SpecVQGAN, despite its high spectral detail, fails to correctly identify video objects, resulting in inaccurate audio generation. While Diff-Foley succeeds in crafting synchronized audio, it occasionally struggles to reproduce the exact sounds associated with specific objects. Overall, our approach demonstrates a superior balance between audio quality, synchronization, and object-specific sound accuracy.
@@
4.2.2. Quantative Results
For each video from VGG-AnimSeg-4k test set, we generate an 8.2-second audio clip for evaluation. Tab. 1 demonstrates our method presents superior performance compared to other methods in terms of FD, FAD, KL, and KID, indicating that our results closely resemble the ground truth. Our excellent CAV score further highlights our method’s exceptional object alignment between the generated audio and input videos.
Figure 7. Comparison of different visual features. Our method (MVE and MVE+CLIP) performs better in capturing the semantic information and temporal conditions in the video, while others generate target-agnostic sounds or low-fidelity audio.
@@ 4.3.1. MVE Ablation
To validate the effectiveness of the visual features extracted by MVE, we compare them with other viusal features. We fine-tune the LDM using different visual features extracted from the training set of VGG-AnimSeg-4k, and then generate 10 audio samples for each test video for evaluation. As shown in Tab. 2, our method achieves best in multiple metrics. Although CAVP features achieve the highest IS score, they perform worse in other metric. Since IS is originally designed to measure the quality of generated images, it does not guarantee the quality of the generated audio. Additionally, adding CLIP features to MVE features (referred to as MVE+CLIP) further enhances performance. This improvement is attributed to the rich and high-level semantic information provided by CLIP features, which may be ignored in convolutional backbone. Fig. 7 shows a qualitative comparison of a representative example.
@@
Table 2. Investigate the effect of conditioning on different visual features. The results showcase the superiority of MVE features, and MVE+CLIP further enhances performance.
@@ 4.3.2. MLM and RVS Ablation
We conduct an ablation study to validate the effectiveness of RVS and MLM. As shown in Tab. 3, MLM significantly improves the scores of FD, FAD, KL, KID, and CAV. The qualitative comparison in Fig. 8 demonstrates that the model trained with MLM generates audio with more pronounced changes in loudness corresponding to the distance of the target object. However, when RVS is added to our base model, there is a slight decrease in performance. This may be due to the resizing process in RVS, which can distort the aspect ratio of the original images. Despite this, Fig. 9 demonstrates that RVS enables the model to handle multi-object scenes more effectively, highlighting its importance in complex scenarios.
@@
Table 3. Validation of MLM and RVS.
Figure 8. An ablation study of MLM. When the donkey runs closer, our results show a clearer increase in loudness, and when it moves away, the sound diminishes.
Figure 9. An ablation study of RVS. RVS enables the model to generate corresponding sounds when selecting multiple objects.

### v2

@@ 4.1.1. Dataset
Our models are trained and evaluated on the VGG-AnimSeg dataset, consisting of 27,200 training samples and 2,720 testing samples. When data augmentation via Random Video Stitching (RVS) is applied, an additional 6,800 augmented training samples are incorporated, resulting in 34,000 samples for model training.
@@ 4.1.2. Evaluation Metrics
For quantitative assessment, we use Frechet Distance (FD), Frechet Audio Distance (FAD), Inception Score (IS), Kullback-Leibler Divergence (KL) and Kernel Inception Distance (KID) following other audio generation studies (Iashin and Rahtu, 2021; Liu et al., 2023, 2024; Luo et al., 2024; Xing et al., 2024). FD, FAD, KL and KID are employed to quantify the similarity between the generated audio and the ground truth, while IS is utilized to assess the diversity and quality of the generated audio samples. To better evaluate the correspondence between video and audio, we introduce a new metric, the CAV score. The score leverages the model from C-MCR (Wang et al., 2023), which integrates CLIP (Radford et al., 2021) and CLAP (Elizalde et al., 2023) to provide audio-image contrastive representations. For each audio-video pair, we obtain per-frame image embeddings and an audio embedding, and then calculate the similarity between the average image embedding and the audio embedding. A higher CAV score indicates a better match between the generated audio and the original video.
@@
Figure 6. Visualization of generated audio samples: (a) Reproduction of ground truth sounds with preserved synchronization, demonstrating accurate capture of transient motion. (b) Audio volume reduction corresponding to cows running away, while maintaining finer details.
@@ 4.1.3. Implementation Details
For the VOS setting, the IoU threshold and Non-Maximum Suppression (NMS) threshold of SAM are set to 0.88 and 0.8. TAM is configured to use 15 voting frames. For data preprocessing, we resample 10-second video clips to 4 fps and resize each frame to 224$\times$224 pixels. For audio samples, we compute Mel spectrograms with a hop size of 250 and 128 mel bins during OCAV training, and a hop size of 256 and 128 mel bins during LDM training. Considering the limited size of our dataset, we construct our framework based on Diff-Foley and fine-tune it using pre-trained weights, since training from scratch is insufficient. Specifically, the $f_{v}$ and $f_{m}$ in MVE utilize the SlowOnly (Feichtenhofer et al., 2019) architecture, known for its proficiency in detecting temporal action changes within videos. $f_{a}$ leverages an audio encoder provided by PANNs (Kong et al., 2020). The architecture of LDM, including the latent encoder and decoder, is inherited from Stable Diffusion-V1.4 (SD-V1.4) (Rombach et al., 2022). During the OCAV training process, we randomly sample 4-second segments from the original audio and videos, denoted as $\mathcal{A}\in\mathbb{R}^{256\times 128}$ and $\mathcal{V}\in\mathbb{R}^{16\times 224\times 224\times 3}$, yielding corresponding features $\boldsymbol{x}_{a}\in\mathbb{R}^{16\times 512}$, $\boldsymbol{x}_{v}\in\mathbb{R}^{16\times 512}$. For LDM training, we first extract visual features through MVE from the 10-second videos, resulting in $\boldsymbol{x}_{v}\in\mathbb{R}^{40\times 512}$. Following the Im2Wav (Sheffer and Adi, 2023) method, we extract per-frame CLIP embeddings, resulting in $\boldsymbol{x}^{*}_{v}\in\mathbb{R}^{40\times 512}$, and add them to $\boldsymbol{x}_{v}$ to form the condition embedding $\boldsymbol{x}_{c}$ for the LDM. At the inference stage, we set the Classifier-Free Guidance Scale to 4.5, the Classifier-Guidance Scale to 50, and use the DPM-Solver (Lu et al., 2022) sampler with 50 inference steps.
@@ 4.2.1. Qualitative Results
We conduct a comparative experiment of our method with several recently published, open-source V2A approaches, including SpecVQGAN (Iashin and Rahtu, 2021), Im2Wav (Sheffer and Adi, 2023), Seeing and Hearing (SAH) (Xing et al., 2024) and Diff-Foley (Luo et al., 2024). The qualitative evaluation focuses on three main aspects: the capacity to produce coherent audio, the relevance of the generated audio to the target objects, and the temporal alignment between audio and visual content, particularly the movements. As shown in Fig. 6, our method excels in generating audio most similar to the original for given video inputs, accurately capturing nuances such dog barks and cow moos, while maintaining synchronization with their movements. Other methods fall short, either producing visually accurate but poorly synchronized or low-quality audio (Im2Wav) or generating irrelevant sounds (SAH). SpecVQGAN, despite its high spectral detail, fails to correctly identify video objects, resulting in inaccurate audio generation. While Diff-Foley succeeds in crafting synchronized audio, it occasionally struggles to reproduce the exact sounds associated with specific objects. Overall, our approach demonstrates a superior balance between audio quality, synchronization, and object-specific sound accuracy.
@@
4.2.2. Quantitative Results
For each video from VGG-AnimSeg-4k test set, we generate an 8.2-second audio clip for evaluation. Tab. 1 demonstrates our method presents superior performance compared to other methods in terms of FD, FAD, KL, and KID, indicating that our results closely resemble the ground truth. Our excellent CAV score further highlights our method’s exceptional alignment between the generated audio and input videos.
Figure 7. Comparison of audio generation results using different visual features, demonstrating that MVE and MVE+CLIP capture semantic information and temporal context more effectively than other methods, which produce target-agnostic or low-fidelity audio.
@@ 4.3.1. MVE Ablation
To validate the effectiveness of the visual features extracted by MVE, we compare them with other visual features. We fine-tune the LDM using different visual features extracted from the VGG-AnimSeg-4k training set, and then generate 10 audio samples for each test video for evaluation. As shown in Tab. 2, our method achieves best in multiple metrics. Although CAVP features achieve the highest IS score, they perform worse on other metrics. Since IS is originally designed to measure the quality of generated images, it does not guarantee the quality of the generated audio. Additionally, adding CLIP features to MVE features (referred to as MVE+CLIP) further enhances performance. This improvement is attributed to the rich, high-level semantic information provided by CLIP features, which a convolutional backbone may ignore. Fig. 7 shows a qualitative comparison of a representative example.
@@
Table 2. Investigation of the impact of conditioning the model on different visual features, demonstrating that MVE features outperform other methods, and MVE+CLIP features lead to further performance improvements.
@@ 4.3.2. MLM and RVS Ablation
We conduct an ablation study to validate the effectiveness of RVS and MLM. As shown in Tab. 3, MLM significantly improves the scores of FD, FAD, KL, KID, and CAV. The qualitative comparison in Fig. 8 demonstrates that the model trained with MLM generates audio with more pronounced changes in loudness corresponding to the distance of the target object. However, adding RVS to our base model slightly decreases performance. This may be due to the resizing process in RVS, which can distort the aspect ratio of the original images. Despite this, Fig. 9 demonstrates that RVS enables the model to handle multi-object scenes more effectively, highlighting its importance in complex scenarios.
@@
Table 3. Validation of the Masked Loudness Modulation (MLM) and Random Video Stitching (RVS) components through quantitative metrics.
Figure 8. Ablation study of the Masked Loudness Modulation (MLM) component, demonstrating that as the donkey moves closer, the generated audio exhibits a clearer increase in loudness, and diminishes as it moves away.
Figure 9. Ablation study of Random Video Stitching (RVS), demonstrating its role in enabling the model to generate distinct sounds corresponding to multiple selected objects in the scene.